Monday, 24 August

Special Activities

\underline{14:00 – 18:00 H 104}

Session on Electronic Publishing Chair: Peter Michor, Vienna, Austria

14:00 – 14:30 Jane Kister, Mathematical Reviews Bernd Wegner, Zentralblatt für Mathematik The year 2000 Mathematics Subject Classification, MSC2000

14:30 – 15:00 Judith Plümer, Universität Osnabrück, Germany Mathematical Metadata

15:00 – 15:30 Stephen Buswell, Stilo Technology, UK An Introduction to MATHML

15:30 – 15:45 Discussion
15:45 – 16:15 Coffee break

16:15 – 18:00 Panel Discussion The future of electronic communication, information and publishing Panelists: N. Calkin, J. Ewing, M. Grötschel, J. Heinze, C. Hutchins, P. Michor, R. Palais Chair: David Mumford, IMU president

ICM’98 – Program  1